OpenStudy (anonymous):

how to integrate this? x^(1/3)/[x^(1/3) -1] dx

OpenStudy (psymon):

let u = x^(1/3)-1 Therefore we can also say x^(1/3) = u+ 1 if we solve for it. Usuing these subs, I can rewrite our problem. \[\int\limits_{}^{}\frac{ u+1 }{ u }=\int\limits_{}^{}1 + \int\limits_{}^{}\frac{ 1 }{ u } \]

OpenStudy (anonymous):

ur da best, thank you ^_____^

OpenStudy (psymon):

Yep, np :3